what is the sum of three consecutive odd integers if the sum of the second and third numbers is six more than two times the first number?
Answer by Earlsdon(6103)

Let (2x-1) be the first odd number, ( x can be any integer), so (2x-1+2 = 2x+1) will be the next cosecutive odd integer and (2x-1+4 = 2x+3) will be the third consecutive odd integer.
From the problem description, we have:
(2x+1)+(2x+3) = 2(2x-1)+6 Simplifying, we have:
4x+4 = 4x+4
Well, as you can see, we have an identity which means that x can be any value.
So the answer to the problem is that you can pick any three consecutive odd integers at random and their sum will be the solution.
For example:
17, 19, and 21
19+21 = 2(17)+6
40 = 34+6
40 = 40